to land, an airplane will aproach an airport at a 3 degree angle of depression. If the plane is flying at 30,000ft. fint the ground distance from the airport to the point directly below the plane when the pilot begins descending. give your answer to the nearest 10,000 ft.

what is the answer?

Answer by stanbon(29603)

to land, an airplane will aproach an airport at a 3 degree angle of depression. If the plane is flying at 30,000ft. find the ground distance from the airport to the point directly below the plane when the pilot begins descending. give your answer to the nearest 10,000 ft.
----------------
Draw the picture.
tan(3 degrees) = [altitude]/[ground distance from airport
[ground distance] = 30,000/tan(3 degrees)
ground distance = 30,000/0.0524077793...=572,434.10 ft.
By the way, this is more than 108 miles.
Cheers,
Stan H.